OpenStudy (anonymous):

If x² - y² = 55, and x - y = 11, then y = A. 8 B. 5 C. 3 D. -8 E. -3

OpenStudy (haseeb96):

E.-3 is the corrrect answer

OpenStudy (anonymous):

E. -3 thats right

OpenStudy (anonymous):

If x-y = 11 then y = x-11 x²-(x-11)² =55 x²-(x²-22x+121) = 55 22x - 121 = 55 22x = 176 x =8 y = x-11 y = 8-11 y = -3. E